12sinx-5coax find the maximum and minimum belie of the question +2 1st year

Grade:11

1 Answers

Arun
6 years ago
We know that in
a sinx - b cosx
Maximum value = \sqrt(a2 + b2)
Minimum value =  - \sqrt(a2 + b2)
hence in this equation
Maximum value = \sqrt(144+25) = 13
Minimum value = - \sqrt(144+25) = -13